OpenStudy (anonymous):

let sin4theta=1/2 and tan4theta>0, find cos2theta

OpenStudy (anonymous):

\[\sin 4\theta = \sin \Pi/6\]

OpenStudy (anonymous):

\[\theta=\Pi/24\]

OpenStudy (anonymous):

book says \[1/2\sqrt{2+\sqrt{3}}\]

OpenStudy (anonymous):

lol sorry misunderstanding..

OpenStudy (anonymous):

we are looking for cos2(theta)

OpenStudy (anonymous):

wait...lol

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

\[\cos2\theta = 1-2\sin^2\]

OpenStudy (anonymous):

use this

OpenStudy (ajprincess):

@Yahoo! \(\theta\) is only \(\large\frac{\pi}{24}\) and not \(\sin \theta\)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

oh..sorryyy

OpenStudy (anonymous):

\[\sin \Pi/24\]

OpenStudy (anonymous):

yw...

OpenStudy (ajprincess):

it's k.:)